OpenStudy (anonymous):

Choose the correct slope of the line that passes through the points (-4, 8) and (-3, -6). (2 points) Question 1 options: 1) 7 2) -14 3) 0 4) 14

OpenStudy (anonymous):

slope=(-6-8)/(-3--4) = -14/1 =-14